LOFAR Newsletters October 2016

Distributed on 24 October

 

Announcements:

--------------------------

 

-       After a very protracted process of consultation, study, and negotiation, the NL Ministry of Economic Affairs has decided to allow the construction of a windmill park in the vicinity of the LOFAR Core, but under strict conditions that provide significant mitigation to LOFAR. It will be a few years before the windmills start to operate. The coming few months will be used to enter into a dialogue with our user community, both to discuss how their particular projects may be impacted, and on how to ensure that LOFAR/LOFAR2.0 can best continue its position as a top-ranked scientific observatory. More details will be given at the LSM on 26 October.

-       Between April 14th 2014 (LOFAR Release 2.2) and September 13th 2016 (LOFAR Release 2.16.7) there was an issue with the LOFAR antenna cable delay compensation not being properly set, which has potentially affected some observations. The issue was caused by RCUs keeping the previous cable delay settings (and ignoring new ones), if the RCUs were still powered on after the end of a previous observation. This happened whenever there was a gap smaller than about 3 minutes between observations and if the switch was from HBA to any LBA mode, or between any two LBA modes. These are not very frequent combinations. The PI's of all affected projects have been informed. More details, including information on how the issue manifests itself in the data, are given here.

-       The Conference ‘The Broad Impact of Low Frequency observing' will take place between 19-23 June 2017 in Bologna (Italy). In November, we will circulate the second announcement and open the registration.

-       32 regular proposals were submitted for Cycle 7, which will start on 15 November 2016. The oversubscription factors for the observing and processing categories are 2.7 and 1.7, respectively. The PC meeting will take place on 7-8 November, at ASTRON.

-       A Junior Telescope Scientist position for LOFAR is available within the Radio Observatory of ASTRON. The advertisement can be found here. Deadline for submission of applications: 1 November 2016.

 

 

Array status:

----------------------

-       38 stations operational in the Netherlands: 24 core stations and 14 remote stations.

-       12 international stations operational: DE601, DE602, DE603, DE604, DE605, FR606, SE607, UK608, DE609, PL610, PL611, PL612

-       A new LOFAR station (IE613) will be built on the grounds of Birr Castle, located centrally in Ireland. Expected completion: first half 2017.

-       Oscillating tiles were detected during BF observations at a few stations and were deactivated.

-       Maintenance of LOFAR stations is in progress and will continue till the weather conditions will allow.

-       The overview of non-operational antenna elements for LBA and HBA is available here.

-       Station calibration:

  • New LBA Outer tables were installed on Aug 1st on the Dutch array. Data for HBA High have been collected and manual flagging of the data needs to be performed. 
  • The software development necessary to perform calibration in LBA Sparse modes will be finalized in the coming weeks.

 

Observing System Status

------------------------------------ 

-       The observing system experienced instability during the period September - October 2016. Problems were mainly associated with the processing and LTA (ingest) systems and COBALT nodes. These caused delays in the exploitation of the Cycle 6 observing programme and in digesting data out of the production cluster(s).

-       The new CEP4 cluster is under commissioning. While some issues still remain (see next section), the commissioning of the imaging and TBB observing modes is completed and that of the Beamformed mode is close to completion. Various production observations have been shifted to the next Cycle to accommodate CEP4 commissioning.

 

 

Software Development Status (J. Annyas):

------------------------------------------

-       CEP4 nodes crashes were experienced during commissioning while increasing the observing & processing load on the system. These were temporarily fixed with a workaround, but a better solution is needed. CEP4 data loss is the next major issue that needs attention. The new web scheduler is well liked for its flexibility and overview. 

-       A new model LCU has been installed on CS302. Tests in production will take place at the end of October. When successful, more stations will be upgraded through a step-by-step approach. We aim at completing this in November, if the weather will allow. 

-       Knowledge transfer of Astrowise/LTA Portal has started with the aim to take support and development into Astron.

-       The responsive telescope project was started up. The backlog is in place in its crude form and is merged with the leftover from the Resource Assigner.

-       A new and more robust ingest code has been developed.

-       New team members are hired to fill the open positions.

 

 

 

CITT Status (E. Orru' & T. J. Dijkema)

--------------------------------------------------------------------

-       The pre-factor pipeline release 2.0 is ready for production. Work is on going in order to allow ingesting the output products into the LTA.

-       FACTOR pipeline: the code was heavily updated to use WSclean for imaging and NDPPP for solving. The new Dysco compression method (A. Offringa) was added in order to limit the size of the data products and the I/O request. The processing time has been drastically reduced making it competitive with respect to other software algorithms.

-       WSclean run with IDG. The run was successful but a bug was identified - work to solve this is on going.

-       An LBA workshop took place in Leiden: various strategies are being investigated to solve the direction dependent calibration at these frequencies. A prototype pipeline for direction independent calibration has been tested and some issues arose. Changes to the recipe and bug solving need to be applied before more extensive testing can take place. More information about this topic will be given at the LSM on 23 November.

 

 

MSSS Status (J. Broderick & G. Heald)

--------------------------------------------------------------------

-       A new MSSS-HBA internal data release (v0.3) is now available on the ASTRON VO server. The updated catalog has a far more reliable flux density scale, particularly in relation to the in-band spectral behaviour. Various aspects of the mosaicking, source finding and catalog production procedures have also been refined.

-       A campaign to reprocess the MSSS-HBA data at higher angular resolution will commence in the coming weeks. In preparation, benchmarking is being performed on CEP3 and other external clusters.

 

Observing Programmes

--------------------------------------------------------------------

-       Cycle 6 observing programme: 67% complete. The observing schedule can be fund here.

-       Cycle 5 observing programme: 92% complete. The rest is being observed with second priority during Cycle 6.

 

 

 

CEP news:

----------------

-       CEP4

  • The CEP4 cluster is being commissioned.

-       CEP2:

  • Disk space situation is often critical, especially in case of instability of the processing and ingest systems and after week-ends.
  • During the past two months, a few locus nodes experienced issues and were temporarily taken out of production.
  • Locus095 has been cannibalized to repair other locus machines.

 

-       CEP3:

  • Cluster info and schedule available here.
  • Monitoring of node usage is active. Statistics will be used to evaluate requests for extensions.

 

 

Documentation:

-------------------------------------------------------------------

Note: we remind the reader that an important source of documentation about LOFAR is available here.

-       The pdf's and video recordings of all the lectures given at the 4th LOFAR Data Processing School are available here.

 

 

Calendar next LOFAR activities:

-------------------------------------------------------------------

Note: the following events are now marked on an online calendar that is available here.

 

-       Next LSM's : 26/10, 23/11

(all presentations given at the LSM and video recordings are available here).

-       ‘The Broad Impact of Low Frequency observing': 19-23 June 2017, Bologna (Italy).

-       Next Stop days: 01/11, 06/12

-       Next software roll outs: 07/11, 12/12

-       Next LOFAR bulletin: December 2016

Design: Kuenst.    Development: Dripl.    © 2020 ASTRON